The scale of mercury thermometer is graduated in milimeters. It reads 106mm and 108.6 mm at two standard fixed points. What reading will this thermometer give  in a liquid of temperature 72°F?????


 


Grade:10

1 Answers

neeraj agarwal
10 years ago
let x is the reading ..... x - 106 / 108.6 - 106 = 72- 32 / 212 - 32 x - 106 / 2.6 = 40 / 180 x - 106 = 2 / 9 * 2.6 x= 106.578